OpenStudy (anonymous):

One number is 7 more than another number. The sum of the numbers is 63. What are the numbers? A. 10 and 7 B. 30 and 37 C. 28 and 35 D. 54 and 9

OpenStudy (anonymous):

x = a number y = the other number y + 7 = x x + y = 63 now sub y + 7 in for x in the 2nd equation

OpenStudy (anonymous):

do you understand ?

OpenStudy (anonymous):

so wait, i do y + 7 = 63 ?

OpenStudy (anonymous):

no....sub in y + 7 in the 2nd equation....like this x + y = 63 (y + 7) + y = 63

OpenStudy (anonymous):

now combine like terms

OpenStudy (anonymous):

okay, so them as a whole equation makes 35 and 28 if i just do the x+y=63 then (y+7) + y = 63

OpenStudy (anonymous):

atleast thats what my calculator shows is 35 and 28

OpenStudy (anonymous):

let me check y + 7 + y = 63 2y + 7 = 63 (subtract 7 from each side) 2y = 63 - 7 2y = 56 y = 28 x + y = 63 x + 28 = 63 x = 63 - 28 x = 32 you got it :)

OpenStudy (anonymous):

i know i could have just done that in the first place, but i needed to know the formula for it so when i go for my tests i have the formula! thank you so much for the explanation!!

OpenStudy (anonymous):

The hardest part about problems like this is the set up. Because if it is not set up correctly, it messes the whole problem up. Glad to be of help :)